# Neural-Quantum-States Impurity Solver for Quantum Embedding Problems

### v1

@@ Abstract
Neural quantum states (NQS) have emerged as a promising approach to solve second-quantised Hamiltonians, because of their scalability and flexibility. In this work, we design and benchmark an NQS impurity solver for the quantum embedding methods, focusing on the ghost Gutzwiller Approximation (gGA) framework. We introduce a graph transformer-based NQS framework able to represent arbitrarily connected impurity orbitals and develop an error control mechanism to stabilise iterative updates throughout the quantum embedding loops. We validate the accuracy of our approach with benchmark gGA calculations of the Anderson Lattice Model, yielding results in excellent agreement with the exact diagonalisation impurity solver. Finally, our analysis of the computational budget reveals the method’s principal bottleneck to be the high-accuracy sampling of physical observables required by the embedding loop, rather than the NQS variational optimisation, directly highlighting the critical need for more efficient inference techniques.
@@ 1.  Introduction
Materials with strong electronic correlations have been the subject of extensive investigation in physics and materials science for decades. Such systems exhibit a variety of electronic phases, including metallic, insulating, and superconducting states [1], and represent a versatile design space for technological applications in electronics, quantum computing, and sensing. Designing new correlated materials with targeted properties therefore depends on the ability to solve the many-body electronic Hamiltonian, a computationally demanding task [2].
@@
Quantum embedding (QE) methods provide a robust framework for overcoming these challenges [3, 4, 5]. The common strategy underlying these methods is to describe each fragment of interacting orbitals through an effective model, where its complex environment is replaced by a simpler, entangled quantum bath, designed to approximate its influence [3, 6]. The link between this effective model and the original system is established through self-consistency conditions, and different embedding schemes are defined by their choice of which physical property to match [3]. Dynamic mean-field theory (DMFT)  [7, 8, 9, 10, 11, 12, 13, 14], for instance, uses frequency-dependent one-body Green’s functions, whereas density matrix embedding theory (DMET) [5, 15, 16, 17, 18, 19] typically uses one- and two-body density matrices. The recently developed gGA [20, 21, 22, 23, 24, 25, 26] is a powerful variational method that generalizes the standard Gutzwiller Approximation (GA) [27, 28, 29, 30, 31, 32, 33, 34] by systematically extending its variational space with auxiliary “ghost” fermionic degrees of freedom. This approach yields results in remarkable agreement with DMFT but at a much lower computational cost, as it requires calculating only the ground state of a finite-size impurity model, whereas DMFT requires the full spectra from an impurity model that corresponds to an infinite bath. Successful applications of gGA include accurate modelling of the Anderson lattice systems [21], excitonic phenomena [20], non-equilibrium systems [35] and altermagnetic systems [36], along with extensions that achieve charge self-consistency with density functional theory (DFT) [22], demonstrating its versatility and practical utility in real-material contexts.
@@
Figure 1: The neural quantum states - quantum embedding (NQS-QE) workflow. The many-body Hamiltonian is embedded into impurity models via quantum embedding algorithms. It is then solved by the NQS impurity solver. The observables are derived from the graph neural network representation of the many-body wave function, which is then used to update the impurity parameters, thereby forming a self-consistent loop.
@@
Despite the advances offered by QE methods, solving the impurity model remains a central computational bottleneck. Current impurity solvers include Exact Diagonalisation (ED) [37, 38], Numerical Renormalisation Group (NRG) [39, 40], Density Matrix Renormalisation Group (DMRG) [41, 42], and Quantum Monte Carlo (QMC) [43, 44]. While each of these methods has distinct advantages, their computational complexity often limits their application to practically relevant systems.
@@
Neural Quantum States (NQS) [45, 46, 47, 48, 49, 50], due to their inherent flexibility in representing wave functions in second-quantised formalisms, offer a compelling alternative to some of these challenges. In particular, their variational foundation provides us with controlled approximation errors through parameter optimisation [51]. Additionally, variational Monte Carlo provides us with a favourable computational scaling [52], especially advantageous for multi-orbital scenarios. The rapidly developing field of NQS shows promising trends toward a unified and scalable impurity solver [53, 54]. Nonetheless, impurity solvers based on NQS must address specific requirements, such as handling irregular hopping patterns that demand arbitrary orbital interactions and ensuring numerical stability and consistency. Accurately calculating observables from the optimised wave function further introduces additional methodological challenges.
@@
In this work, we develop a graph transformer-based NQS impurity solver to represent arbitrarily connected impurity orbitals. We implement a systematic error control framework to rigorously bound uncertainties arising during the iterative optimisation and embedded properties sampling of the NQS wave function. Our model initialises node features with spin configurations and employs iterative graph attention mechanisms, achieving high flexibility in wave function representation. Optimisation efficiency is enhanced using recently proposed algorithms such as the Neural Network Pfaffian [55], MinSR[47] and SPRING[49]. Coupling our graph-based NQS solver with the gGA quantum embedding method, we validate its performance using benchmark Anderson Lattice Models [21], obtaining excellent agreement with ED solutions. Our results demonstrate the feasibility of employing NQS as impurity solvers within quantum embedding algorithms. Furthermore, we conduct extensive analyses of convergence behaviours across different error control parameters and computational timing, highlighting critical open questions that must be addressed to achieve scalable and practically relevant NQS-based impurity solvers.
@@
The rest of the paper is organised as follows. In the next section, the technical details of the method are presented, including the NQS impurity solver, its associated error control, as well as its integration with the ghost Gutzwiller quantum embedding approach. In section 3, benchmark results associated with the Anderson Lattice Model and a detailed cost and error analysis of the NQS solver are illustrated. Finally, section 4 presents a discussion and a short summary of the NQS-gGA many-body formalism developed in this work.
@@
Figure 2: (a) The workflow for the neural quantum states (NQS) impurity solver; The outer loop shows the impurity solver’s coupling with the quantum embedding method, which generates an impurity model for the NQS to solve. The solving process includes the NQS optimisation and properties sampling steps, with error control metrics E-tol and P-tol to guarantee convergence. The metrics are checked with tolerance $\epsilon_{t}$ and $\epsilon_{p}$ for energy and properties, respectively. Both steps are driven by the Monte Carlo (MC) sampler to compute expectations. (b) The neural network wave function’s architecture. The spin orbital configuration is first encoded with the occupation and positional embedding method, and passes through the main body of the network, which comprises N blocks of graph-transformer-like architecture with a residual connected graph attention (GAT) layer and a feedforward neural network (FFN) layer.
@@ 2.  The NQS-gGA formalism
In this section, we present the technical details of the NQS-gGA formalism developed in this work. As illustrated in Fig.2 (a), the impurity model, generated by the quantum embedding outer loops, will be fed into the NQS solving process, which consists of the NQS optimisation and properties sampling. For each step in the solving process, we have an error control metric to ensure the variance is minimised. In the following subsections, we will discuss the NQS method we used to construct the impurity solver in part A, the error control system that guarantees the stability of solving iterations in part B. In part C we will outline the gGA algorithmic structure.
@@ II.1 Neural Quantum States Impurity Solver
The neural quantum state acts as a many-body wave function ansatz in the variational quantum Monte Carlo framework. Specifically, we consider an NQS wave function, expressed in second quantisation as follows:
@@
where $\mathbb{E}_{x\sim\Psi_{\theta}^{2}}$ represents the expectation over the probability distribution of the current parameterized wave function $\Psi_{\theta}$. We employ the Markov chain Monte Carlo [57] to perform sampling and the neighbour-exchange method [56] to generate new electron configurations. To approach the ground state, the NQS is optimised using stochastic reconfiguration [58] based on a recently proposed MinSR [47] and the SPRING [49] methods to stabilise and accelerate the calculation.
@@
More specifically, our graph-based model treats the orbital as graph nodes and single-particle interactions as edges. As illustrated in Fig.2 (b), we initialise the node message as the orbital occupation number vector, which is a binary vector of value 1/0. Then, the state vectors are passed into the occupation positional encoding layer, where four different occupations in one site ($|11\rangle,|01\rangle,|10\rangle,|00\rangle$) are mapped to different vector embeddings, as is commonly adopted in natural language processing [62]. Meanwhile, the positional information is added to the embedding vectors via the positional embeddings [62] to uniquely distinguish each orbital site, which results in our final embeddings with all necessary state information. Furthermore, the embeddings are passed through the main body of the model, backboned by the GATv2 [63] graph attentional architecture. We mimic the transformer architecture by composing each iterative layer using one graph attention (GAT) layer and a feedforward neural network (FFN) layer with residual connections. After iterative accumulation, the features are then mapped via a linear layer to two real scalar quantities, corresponding to the wavefunction amplitude and the complex phases, as our wave function’s output $\langle x|\Psi_{\theta}\rangle$
A significant challenge in building a stable solver is to define a clear error metric for the wave function optimisation. Conventional VMC provides an intuitive guideline — that lower variance typically correlates with reduced error, but the relationship is not quantitative. We will discuss this important issue in the following section.
@@ II.2 Error Control System
A characteristic functional distinction between solving standalone Hamiltonians and acting as solvers to form self-consistent solutions is the requirement for numerical stability. In quantum embedding frameworks, the impurity model and the physical system are self-consistently coupled. The solution to the impurity model directly influences the state of the original system, which, in turn, affects the updates of the impurity problem itself. Uncontrolled errors can propagate through iterative loops, causing instability in the embedding procedure. Consequently, establishing a systematic error control methodology for the NQS solver within the variational Monte Carlo (VMC) framework is essential.
The errors arising in the VMC-NQS process can be categorised into two primary sources: (1) the wave function optimisation error; (2) the Monte Carlo sampling error. The wave function optimisation error refers to the discrepancy between the converged NQS wave function and the true ground-state wave function, which appears during the NQS optimisation steps in Fig.2 (a).
The quality of the optimised NQS wave function, $|\Psi_{\theta}\rangle$, is determined by how closely it approximates a true eigenstate of the impurity Hamiltonian. The most direct measure of this is the energy variance, $\text{Var}[E]$, which must vanish for an exact eigenstate and is therefore the fundamental diagnostic monitored during optimisation:
@@
which is a more robust, dimensionless, and size-independent metric of variational accuracy. Here $N$ is the number of sites, and the denominator provides a normalising energy scale, where $E$ is the variational energy and $E_{T}$ is a fixed reference energy.
We obtain this reference energy by averaging the energies $\langle\Psi_{\epsilon}|\hat{H}|\Psi_{\epsilon}\rangle$ of several randomly-generated slater determinants, $|\Psi_{\epsilon}\rangle$, each of which can be evaluated analytically using the Hartree-Fock formula at a negligible computational cost. This method avoid the explicit Monte Carlo sampling used in [64] when treating $|\Psi_{\epsilon}\rangle$ as a true random many-body states
The second major source of error is the statistical uncertainty from the Monte Carlo sampling performed during both the NQS optimisation and properties sampling steps (see Fig. 2(a)). The numerical expectation of the observable from the Monte Carlo sampling is distributed as a Gaussian variable, whose standard deviation scales inversely as the square root of the sample size $M$ [56], i.e. $\sigma\sim O(1/\sqrt{M})$. This brings two critical uncertainties in the NQS solver:
@@
    Firstly, when using the V-score as the wave function optimisation metric, it carries intrinsic statistical uncertainty due to finite sampling. This uncertainty is addressed by applying a confidence interval correction proportional to $O(1/\sqrt{M})$ to the computed V-score; we denote the arrived metric score as E-tol.
@@
    Secondly, when sampling embedded physical properties from the optimised wave function, it introduces direct uncertainty into the quantum embedding loop. This uncertainty must be independently controlled, adhering strictly to the convergence criteria of the outer embedding iterations (see big blue arrows in Fig.1). To bound this error, we apply a stringent criterion, typically $3\sigma$, corresponding to three times the standard deviation of the sampling expectation. This metric should be set lower than the convergence requirements of the embedding method. We denote the arrived properties sampling error metric as P-tol.
Notably, this convergence of P-tol often dominates the computational cost. For example, to reach a 0.001 residual for the embedding algorithm, it requires the error bound $3\sigma$ to be smaller than $5\times 10^{-4}$, corresponding to a standard deviation of approximately $\sigma=1.7\times 10^{-4}$. This precision necessitates sampling in the order of $10^{8}$ configurations (as $M\propto 1/\sigma^{2}$), often exceeding the computational resources required for the wave function optimisation itself.
By integrating the P-tol and E-tol metrics, we systematically control both the wave function optimisation and sampling errors. This guarantees the convergence stability of the NQS impurity solver. We will demonstrate the practical effectiveness of this combined error-control strategy in Section 3 below.

### v2

@@ Abstract
Neural quantum states (NQS) have emerged as a promising approach to solve second-quantized Hamiltonians, because of their scalability and flexibility. In this work, we design and benchmark an NQS impurity solver for the quantum embedding (QE) methods, focusing on the ghost Gutzwiller Approximation (gGA) framework. We introduce a graph transformer-based NQS framework able to represent arbitrarily connected impurity orbitals of the embedding Hamiltonian (EH) and develop an error control mechanism to stabilize iterative updates throughout the QE loops. We validate the accuracy of our approach with benchmark gGA calculations of the Anderson Lattice Model, yielding results in excellent agreement with the exact diagonalisation impurity solver. Finally, our analysis of the computational budget reveals the method’s principal bottleneck to be the high-accuracy sampling of physical observables required by the embedding loop, rather than the NQS variational optimization, directly highlighting the critical need for more efficient inference techniques.
@@ 1.  Introduction
Materials with strong electronic correlations have been the subject of extensive investigation in physics and materials science for decades. Such systems exhibit a variety of electronic phases, including metallic, insulating, and superconducting states [49], and represent a versatile design space for technological applications in electronics, quantum computing, and sensing. Designing new correlated materials with targeted properties therefore depends on the ability to solve the many-body electronic Hamiltonian, a computationally demanding task [1]. QE methods provide a robust framework for overcoming these challenges [60, 47, 32]. The common strategy underlying these methods is to describe each fragment of interacting orbitals through an effective EH, where its complex environment is replaced by a simpler, entangled quantum bath, designed to approximate its influence [60, 50]. The link between this effective model and the original system is established through self-consistency conditions, and different embedding schemes are defined by their choice of which physical property to match [60]. Dynamic mean-field theory (DMFT)  [3, 2, 28, 22, 70, 33, 54, 52], for instance, uses frequency-dependent one-body Green’s functions, whereas density matrix embedding theory (DMET) [32, 31, 29, 66, 56, 19] typically uses one- and two-body density matrices. The recently developed gGA [36, 21, 43, 39, 38, 41, 42] is a powerful variational method that generalizes the standard Gutzwiller Approximation (GA) [27, 37, 34, 20, 30, 35, 10, 4] by systematically extending its variational space with auxiliary “ghost” fermionic degrees of freedom. This approach yields results in remarkable agreement with DMFT but at a much lower computational cost, as it requires calculating only the ground state of a finite-size impurity model, whereas DMFT requires the full spectra from an impurity model that corresponds to an infinite bath. Successful applications of gGA include accurate modeling of the Anderson lattice systems [21], excitonic phenomena [36], non-equilibrium systems [25] and altermagnetic systems [23], along with extensions that achieve charge self-consistency with density functional theory (DFT) [43], demonstrating its versatility and practical utility in real-material contexts.
@@
Figure 1: The neural quantum states - quantum embedding (NQS-QE) workflow. The many-body Hamiltonian is embedded into impurity models via QE algorithms. It is then solved by the NQS impurity solver. The observables are derived from the graph neural network representation of the many-body wave function, which is then used to update the impurity parameters, thereby forming a self-consistent loop.
@@
Despite the advances offered by QE methods, solving the embedding impurity Hamiltonian model remains a central computational bottleneck. Current impurity solvers used with QE include exact diagonalization (ED) [11, 13], Numerical Renormalization Group (NRG) [65, 9], Density Matrix Renormalization Group (DMRG) [64, 55], Natural Orbital Renormalization Group[62, 63], Tensor Network [18, 17] and Quantum Monte Carlo (QMC) [26, 53]. While each of these methods has distinct advantages, their computational complexity often limits their application to practically relevant systems.
@@
Neural Quantum States (NQS) [14, 58, 15, 51, 24, 44, 69], due to their inherent flexibility in representing wave functions in second-quantized formalisms, offer a compelling alternative to some of these challenges. In particular, their variational foundation provides us with controlled approximation errors through parameter optimization [40]. Additionally, variational Monte Carlo provides a favorable computational scaling [68] especially for multi-orbital scenarios. Specifically in this work, we adopted the recently proposed hidden pfaffian ansatz, which has a proven scaling of $O(N^{2})\sim O(N^{3})$ [16] with N the number of sites. The rapidly developing field of NQS shows promising trends toward a unified and scalable impurity solver [12]. Recent works [48] incorporated the QiankunNet NQS method with DMET and applied successfully to calculate the ground state properties of atomic systems. Nonetheless, impurity solvers based on NQS must address specific requirements such as handling irregular hopping patterns that demand arbitrary orbital interactions and ensuring numerical stability and consistency, which are required for a robust solver embedded in the QE self-consistency workflow. Accurately calculating observables from the optimized wave functions further introduces additional methodological challenges.
@@
In this work, we develop a graph transformer-based NQS impurity solver for the quantum EH, which is able to represent arbitrarily connected impurity orbitals. We implement a systematic error control framework to rigorously bound uncertainties arising during the iterative optimization and embedded properties sampling of the NQS wave function. Our model initializes node features with spin configurations and employs iterative graph attention mechanisms, achieving high flexibility in wave function representation. Optimization efficiency is enhanced using recently proposed algorithms such as the Neural Network Pfaffian [16], MinSR[15] and SPRING[24]. Coupling our graph-based NQS solver with the gGA method, we validate its performance using benchmark Anderson Lattice Models [21], obtaining excellent agreement with ED solutions. Our results demonstrate the feasibility of employing NQS as an impurity solvers within QE algorithms. Furthermore, we conduct extensive analyses of convergence behaviors across different error control parameters and computational timing, highlighting critical open questions that must be addressed to achieve scalable and practically relevant NQS-based impurity solvers.
@@
The rest of the paper is organized as follows. In the next section, the technical details of the method are presented, including the NQS impurity solver, its associated error control, and its integration with the gGA QE approach. In section 3, benchmark results associated with the Anderson Lattice Model and a detailed cost and error analysis of the NQS solver are illustrated. Finally, section 4 presents a discussion and a summary of the NQS-gGA many-body formalism developed in this work.
@@
Figure 2: (a) The workflow for the neural quantum states (NQS) impurity solver. The outer loop shows the impurity solver’s coupling with the QE method, which generates an impurity model for the NQS to solve. The solving process includes the NQS optimization and properties sampling steps, with error control metrics E-tol and P-tol to guarantee convergence. The E-tol and P-tol criteria are checked against user-defined tolerances $\epsilon_{t}$ and $\epsilon_{p}$ for the optimization and sampling steps, respectively. Both steps are driven by the Monte Carlo (MC) sampler to compute expectations. (b) The neural network wave function’s architecture. The spin orbital configuration is first encoded with the occupation and positional embedding method, and passes through the main body of the network, which comprises N blocks of graph-transformer-like architecture with a residual connected graph attention (GAT) layer and a feedforward neural network (FFN) layer.
@@ 2.  The NQS-gGA formalism
In this section, we present the technical details of the NQS-gGA formalism developed in this work. As illustrated in Fig.2 (a), the impurity model, generated by the QE outer loops, will be fed into the NQS solving process, which consists of the NQS optimization and properties sampling. For each step in the solving process, we have an error control metric to ensure the variance is minimized. In the following subsections, we will discuss the NQS method we used to construct the impurity solver in part A, the error control system that guarantees the stability of solving iterations in part B. In part C we will outline the gGA algorithmic structure.
@@ II.1 Neural Quantum States Impurity Solver
The neural quantum state acts as a many-body wave function ansatz in the variational quantum Monte Carlo framework. Specifically, we consider an NQS wave function, expressed in second quantization as follows:
@@
where $\mathbb{E}_{x\sim\Psi_{\theta}^{2}}$ represents the expectation over the probability distribution of the current parameterized wave function $\Psi_{\theta}$. We employ the Markov chain Monte Carlo [8] to perform sampling and the neighbor-exchange method [6] to generate new electron configurations. To approach the ground state, the NQS is optimized using stochastic reconfiguration [59] based on a recently proposed MinSR [15] and the SPRING [24] methods to stabilise and accelerate the calculation.
@@
More specifically, our graph-based model treats the orbital as graph nodes and single-particle interactions as edges. As illustrated in Fig.2 (b), we initialise the node message as the orbital occupation number vector, which is a binary vector of value 1/0. Then, the state vectors are passed into the occupation positional encoding layer, where four different occupations in one site ($|11\rangle,|01\rangle,|10\rangle,|00\rangle$) are mapped to different vector embeddings, as is commonly adopted in natural language processing [61]. Meanwhile, the positional information is added to the embedding vectors via the positional embeddings [61] to uniquely distinguish each orbital site, which results in our final embeddings with all necessary state information. Furthermore, the embeddings are passed through the main body of the model, backboned by the GATv2 [7] graph attention architecture. We mimic the transformer architecture by composing each iterative layer using one graph attention (GAT) layer and a feedforward neural network (FFN) layer with residual connections. After iterative accumulation, the features are then mapped via a linear layer to two real scalar quantities, corresponding to the wave function amplitude and the complex phases, as our wave function’s output $\langle x|\Psi_{\theta}\rangle.$
A significant challenge in building a stable solver is to define a clear error metric for the wave function optimization. Conventional VMC provides an intuitive guideline — that lower variance typically correlates with reduced error, but the relationship is not quantitative. We will discuss this important issue in the following section.
@@ II.2 Error Control System
A characteristic functional distinction between solving standalone Hamiltonians and acting as solvers to form self-consistent solutions is the requirement for numerical stability. In QE frameworks, the impurity model and the physical system are self-consistently coupled. The solution to the impurity model directly influences the state of the original system, which, in turn, affects the updates of the impurity problem itself. Uncontrolled errors can propagate through iterative loops, causing instability in the embedding procedure. Consequently, establishing a systematic error control methodology for the NQS solver within the variational Monte Carlo (VMC) framework is essential.
The errors arising in the VMC-NQS process can be categorized into two primary sources: (1) the wave function optimization error; (2) the Monte Carlo sampling error. The wave function optimization error refers to the discrepancy between the converged NQS wave function and the true ground-state wave function, which appears during the NQS optimization steps in Fig.2 (a).
The quality of the optimized NQS wave function, $|\Psi_{\theta}\rangle$, is determined by how closely it approximates a true eigenstate of the impurity Hamiltonian. The most direct measure of this is the energy variance, $\text{Var}[E]$, which must vanish for an exact eigenstate and is therefore the fundamental diagnostic monitored during optimization:
@@
which is a more robust, dimensionless, and size-independent metric of variational accuracy. Here $N$ is the number of sites, and the denominator provides a normalizing energy scale, where $E$ is the variational energy and $E_{T}$ is a fixed reference energy.
We obtain this reference energy by averaging the energies $\langle\Psi_{\epsilon}|\hat{H}|\Psi_{\epsilon}\rangle$ of several randomly-generated Slater determinants, $|\Psi_{\epsilon}\rangle$, each of which can be evaluated analytically using the Hartree-Fock formula at a negligible computational cost. This method avoids the explicit Monte Carlo sampling used in [67] when treating $|\Psi_{\epsilon}\rangle$ as a true random many-body state.
The second major source of error is the statistical uncertainty from the Monte Carlo sampling performed during both the NQS optimization and properties sampling steps (see Fig. 2(a)). The numerical expectation of the observable from the Monte Carlo sampling is distributed as a Gaussian variable, whose standard deviation scales inversely as the square root of the sample size $M$ [6], i.e. $\sigma\sim O(1/\sqrt{M})$. This brings two critical uncertainties in the NQS solver:
@@
    Firstly, when using the V-score as the wave function optimization metric, it carries intrinsic statistical uncertainty due to finite sampling. This uncertainty is addressed by applying a confidence interval correction proportional to $O(1/\sqrt{M})$ to the computed V-score; we denote the arrived metric score as E-tol.
@@
    Secondly, when sampling embedded physical properties from the optimized wave function, it introduces direct uncertainty into the QE loop. This uncertainty must be independently controlled, adhering strictly to the convergence criteria of the outer embedding iterations (see big blue arrows in Fig.1). To bound this error, we apply a stringent criterion, typically $3\sigma$, corresponding to three times the standard deviation of the sampling expectation. This metric should be set lower than the convergence requirements of the embedding method. We denote the arrived properties sampling error metric as P-tol.
Notably, this convergence of P-tol often dominates the computational cost. For example, to reach a 0.001 residual for the embedding algorithm, it requires the error bound $3\sigma$ to be smaller than $5\times 10^{-4}$, corresponding to a standard deviation of approximately $\sigma=1.7\times 10^{-4}$. This precision necessitates sampling in the order of $10^{8}$ configurations (as $M\propto 1/\sigma^{2}$), often exceeding the computational resources required for the wave function optimization itself.
By integrating the P-tol and E-tol metrics, we systematically control both the wave function optimization and sampling errors. This guarantees the convergence stability of the NQS impurity solver. We will demonstrate the practical effectiveness of this combined error-control strategy in Section 3 below.